# From Spatial to Actions: Grounding Vision-Language- Action Model in Spatial Foundation Priors

### v1

@@ 7 Hyper-Parameters and Training Details
Real-World Tasks. For real-world evaluation, we initialize the model with pre-trained weights from Kosmos-VLA-2D (OXE) and ESM. Stage 1 training uses a learning rate of 1e-4, a global batch size of 512, and no warmup. Stage 2 hyper-parameters are provided in Tab. 6. We evaluate all VLAs under the following training settings: (1) Base Tasks: models trained separately on three scenarios. (2) Few-shot Adaptation: multi-task training across all four tasks. (3) Spatial-Prompts: efficient fine-tuning on two tasks. Baseline models (SpatialVLA [29], RoboVLM [17]) use the same hyper-parameters as FALCON. For OpenVLA-OFT [14] (LoRA-only), we use a learning rate of 5e-4, global batch size 128, LoRA rank 32, chunk size 5, and Base Tasks train for 150k iterations, other two settings for 100k iterations. All VLAs trained on real-world datasets use only the side camera input.
Across all experiments, both training stages for FALCON use the same number of epochs/iterations, a Constant learning rate scheduler, and the AdamW optimizer. All input images (including depth maps) are resized to a resolution of 224×224
@@ 8 Implementation Details
As demonstrated in Tab. 6, on the CALVIN benchmark, the VLM receives side and wrist camera images with a history length of 16 frames, while the ESM processes third-view images also with same length historical context. Predictions are made using an LSTM-based action predictor that outputs a chunk of $C=10$ future actions. For SimplerEnv and the real-world benchmark, both the VLM and ESM receive a single-step third-view image, and an MLP-based action predictor predicts an action chunk of length $C=5$. During inference, we employ two different action execution strategies: for CALVIN and SimplerEnv, the policy executes the ensemble actions before generating the next chunk. In the real-world setup, the entire action chunk is executed at once. FALCON requires 12.8 GB of GPU memory and runs at approximately 57 Hz on a single NVIDIA RTX 4090 GPU during real-world evaluation.
@@
CALVIN Performance (Tab. 1) LSTM 16 10 Side+Wrist Side 128
2e-5 (ABC→D)
5e-5 (ABCD→D)
5 Ep
@@
Real-World Performance (Fig. 4-7) MLP 1 5 Side Side 512 2e-5 30 Ep
CALVIN Ablation (Tab. 4, Tab. 7) MLP 1 10 Side+Wrist Side 128
5e-5 (ABC→D)
2e-5 (ABCD→D)
5 Ep
@@ 9 Benchmark Details
The physical setup consists of an xArm 6 robot arm equipped with a Robotiq parallel gripper and an Intel RealSense D435i depth camera positioned approximately 0.6 meters away to provide a third-person view, as illustrated in Fig. 9. All fine-tuning datasets are collected via human teleoperation using a Spacemouse device, sampled at 10Hz. We use absolute Cartesian control as the action space for policy training and deployment.
@@ 10 Ablation Study
In this section, we perform ablation studies on CALVIN benchmark to assess key design choices in FALCON, focusing on the impact of spatial token injection positions and modality fusion strategies on its performance.
Table 7: Ablation studies on spatial token injection strategies and fusion methods. Results confirm that the standard FALCON paradigm achieves the best performance, validating it as the optimal design choice.
@@
[…]
How to fusion spatial tokens? We evaluate three modality fusion approaches detailed in Sec. 3.3: cross-attention, FiLM-gated modulation, and element-wise addition (our standard FALCON). As shown in Tab. 7, element-wise addition consistently delivers the best performance across all experimental settings. This method achieves the highest task success rates while remaining both simple and computationally efficient, as it introduces no additional parameters. The results underscore the effectiveness of a straightforward, parameter-free fusion strategy for seamlessly integrating spatial and semantic representations in VLA models.
11 Potential Future Works

### v2

@@ 7 Hyper-Parameters and Training Details
Real-World Tasks. For real-world evaluation, we initialize the model with pre-trained weights from Kosmos-VLA-2D (OXE) and ESM. Stage 1 training uses a learning rate of 1e-4, a global batch size of 512, and no warmup. Stage 2 hyper-parameters are provided in Tab. 7. We evaluate all VLAs under the following training settings: (1) Base Tasks: models trained separately on three scenarios. (2) Few-shot Adaptation: multi-task training across all four tasks. (3) Spatial-Prompts: efficient fine-tuning on two tasks. Baseline models (SpatialVLA [31], RoboVLM [18]) use the same hyper-parameters as FALCON. For OpenVLA-OFT [15] (LoRA-only), we use a learning rate of 5e-4, LoRA rank 32, chunk size 5, and Base Tasks train for 150k iterations, other two settings for 100k iterations. All VLAs trained on real-world datasets use only the side camera input.
Across all experiments, both training stages for FALCON use the same number of epochs/iterations, a Constant learning rate scheduler, and the AdamW optimizer. All input images (including depth maps) are resized to a resolution of 224×224. Besides, we follow [18] to set the loss weighting factor $\lambda=0.01$ for fair comparisons.
@@ 8 Implementation Details
As demonstrated in Tab. 7, on the CALVIN benchmark, the VLM receives side and wrist camera images with a history length of 16 frames, while the ESM processes third-view images also with same length historical context. Predictions are made using an LSTM-based action predictor that outputs a chunk of $C=10$ future actions. For SimplerEnv and the real-world benchmark, both the VLM and ESM receive a single-step third-view image, and an MLP-based action predictor predicts an action chunk of length $C=5$. During inference, we employ two different action execution strategies: for CALVIN and SimplerEnv, the policy executes the ensemble actions before generating the next chunk. In the real-world setup, the entire action chunk is executed at once. FALCON requires $\sim$12.8 GB of GPU memory and runs at approximately 57 Hz on a single NVIDIA RTX 4090 GPU during real-world evaluation.
@@
CALVIN Performance (Tab. 1) LSTM 16 10 Side+Wrist Side 128 2e-5 (ABC→D) 5e-5 (ABCD→D) 5 Ep
@@
Real-World Performance (Fig. 3-5, Fig. 7) MLP 1 5 Side Side 512 2e-5 30 Ep
CALVIN Ablation (Tab. 4-5, Tab. 8, Tab. 10) MLP 1 10 Side+Wrist Side 128 5e-5 (ABC→D) 2e-5 (ABCD→D) 5 Ep
@@ 9 Benchmark Details
The physical setup consists of an xArm 6 robot arm equipped with a Robotiq parallel gripper and an Intel RealSense D435i depth camera positioned approximately 0.6 meters away to provide a third-person view, as illustrated in Fig. 9. All fine-tuning datasets are collected via human teleoperation using a Spacemouse device, sampled at 10Hz. We use absolute Cartesian control as the action space for policy training and deployment [24].
@@ 10 Ablation Study
10.1 Loss Weighting Factors
Table 8: Performance comparison of different loss weighting factors on CALVIN benchmark.
Method Task Tasks Completed in a Row (%) Avg. Len. $\uparrow$
1 2 3 4 5
[…]
Figure 10: Loss curves of FALCON under different weighting factors during training Stage 2 on CALVIN benchmark.
For all experiments in our paper, we set $\lambda=0.01$ for fair comparisons. To further ablate about the sensitivity of different $\lambda$, we conducted an additional experiment for FALCON with $\lambda=0.05$ under RGB-only input on CALVIN zero-shot setting. As shown in Tab. 10.1, performance decreased slightly when $\lambda$ increased from 0.01 to 0.05 (Avg. Len. dropped from 3.91 to 3.87). Analysis of the overall train loss curves (Fig. 10) reveals that larger $\lambda$ leads to increased oscillation during training, while the overall convergence trend remains the same, which indicates that the model is not highly sensitive to small variations in $\lambda$.
10.2 LSTM Head Parameters
Table 9: Performance comparison of Kosmos-VLA (LSTM) under different parameters on CALVIN benchmark.
@@ 10.2 LSTM Head Parameters
Figure 11: Loss curves of Kosmos-VLA (LSTM) under different parameters on CALVIN benchmark.
To systematically evaluate the effects of history length (H) and chunk size (C), we conducted experiments with Kosmos-VLA (RGB-only) using an LSTM action predictor under the CALVIN ABC→D setting.
Performance Comparison. We tested three configurations: (a) H=16, C=10, (b) H=16, C=5, and (c) H=8, C=10. The results shown in Tab. 10.2 demonstrate that H=16, C=10 achieves the strongest performance (Avg. Len. 4.23). Reducing either parameter degrades model performance. Specifically, a shorter history length (H=8) caused the most significant performance drop (Avg. Len. decreased from 4.23 to 3.90), underscoring that richer historical context is crucial for long-horizon tasks in robotic manipulation.
Training Stability. The overall train loss curves (Fig. 11) reveal that while all configurations eventually converge, reducing history length significantly impacts initial training stability. Specifically, H=8, C=10 shows slower descent and greater oscillation in early stages (0-15k steps), whereas reducing chunk size (C from 10 to 5) maintains stable convergence. This indicates that sufficient history length is vital not only for final performance but also for stable optimization.
11 Potential Future Works